OpenStudy (anonymous):

Solve the equation Quantity x minus 1 all over 9 = 2 for x.

mathslover (mathslover):

Can you write an expression for that?

mathslover (mathslover):

x minus 1 can be written as : \( x - 1\) x minus 1 "all over 9" can be written as : \(\cfrac{x-1}{9}\) Right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

So now, it is given that : x minus 1 all over 9 = 2 That is : \(\bf{\cfrac{x-1}{9} = 2}\) Now can you solve for x?

OpenStudy (anonymous):

no

mathslover (mathslover):

Okay, multiply both sides by 9 : \(\cfrac{x-1}{9} \times 9 = 2 \times 9\) \(\cfrac{x-1}{\cancel{9}} \times \cancel{9} = 2 \times 9 \)

mathslover (mathslover):

or , x- 1= 18